Find the number of ways to purchase 5 different kinds of drinks from a selection of 13 drinks

Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(13,5) = (13!)/(5!*(13-5)!)

C(13,5) = (13!)/(5!*8!)

C(13,5) = (13*12*11*10*9*8!)/(5!*8!)

C(13,5) = (13*12*11*10*9)/(5!)

C(13,5) = (13*12*11*10*9)/(5*4*3*2*1)

C(13,5) = (154440)/(120)

C(13,5) = 1287

So there are 1287 ways if order doesn't matter.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------